OpenStudy (anonymous):

The perimeter of a rectangle is 276 inches. The length exceeds the width by 58 inches. Find the length and the width

OpenStudy (anonymous):

What is the equation for the perimeter of a rectangle?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

No. I'm asking you. I'm assuming this must have been introduced to you otherwise they would have stated it explicitly.

OpenStudy (anonymous):

So what is the equation for the perimeter of a rectangle?

OpenStudy (anonymous):

276

OpenStudy (anonymous):

you is confusing me

OpenStudy (anonymous):

No. In general terms. The perimeter of a rectangle = 2*lenght + 2*width

OpenStudy (anonymous):

116

OpenStudy (anonymous):

does u add 116 twice

OpenStudy (radar):

Let x = width then, length would be represented as width + 58 inches or x+58 = length. Now plug this into the equation for perimeter getting 2x + 2(x+58) = 276 2x +2x + 116 = 276 4x = 276 -116 =160 x=160/4=40 width x+58=98 length.

OpenStudy (anonymous):

so is 40 the width and 98 the length

OpenStudy (anonymous):

yes luv